OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12 I really really need help. I'm having a lot of trouble with math.

rishavraj (rishavraj):

write the equation in the form y = mx + c m -- slope c -- y intercept

OpenStudy (anonymous):

I don't understand this at all. It's my worst subject and so far I haven't understood a lot

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@SolomonZelman anyone? I really need help

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@maddiegirl @YungKhris

OpenStudy (anonymous):

whats ur answers?! & have u tried googling this?

OpenStudy (anonymous):

I have and i've tried looking in my textbook and everything and I don't understand. A. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3 B. y = 11/9x - 4/3 slope: 11/9 y-intercept: 4/3 C. y = 11/9x + 4/3 slope: 4/3 y-intercept: 11/9 D. y = -11/9x + 4/3 slope: 11/9 y-intercept: 4/3

OpenStudy (anonymous):

Didn't understand at all. I saw that earlier today

OpenStudy (anonymous):

ok soooo the equation is y=mx+b..........

OpenStudy (anonymous):

i have no idea what that means to be completely honest

OpenStudy (anonymous):

omg this was like freshman year for me like 2 years ago, its guna b hard to remember

OpenStudy (anonymous):

i'm in 12th grade >.< this is algebra 2

OpenStudy (anonymous):

omg im dead ima freaking junior XD

OpenStudy (anonymous):

i thot this was algebra one

OpenStudy (anonymous):

Well algebra 2 is supposed to be in 11th grade. I had to do pre-algebra in 9th grade because my transcripts didn't transfer so that's why i'm kinda behind >.<

OpenStudy (anonymous):

ohhh ok lol well im supposed to b in algebra but i failed geometry so im retaking it

OpenStudy (anonymous):

if i fail this i gotta redo algebra 2 and i gotta wait till next year to start college D:

OpenStudy (anonymous):

well i will help to the best of my knowledge

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome :) soooo... lets c ima have to work this out

rishavraj (rishavraj):

@Emi17 done with this one????

OpenStudy (anonymous):

I don't understand I haven't been able to do it

rishavraj (rishavraj):

slope intercept form is y = mx + c where m is the slope and c is the y-intercept u need to write the given equation in this slope intercept form

OpenStudy (anonymous):

i thot it was mx+b not mx+c

OpenStudy (anonymous):

so... is y=9?... I'm trying my best here

OpenStudy (anonymous):

wait im tryna remember the equation for this

rishavraj (rishavraj):

so it would be hmmmm, \[9y = 11x - 12\] \[y = \frac{ 11 }{ 9 }x - \frac{ 12 }{ 9 }\] \[y = \frac{ 11 }{ 9 }x - \frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

ther u go, now i remember how to work these equations XD

rishavraj (rishavraj):

@Emi17 done ????????

OpenStudy (anonymous):

its B